ASZ8(1L) ASZ8(1L) NAME asz8−crossassemblersformicrocomputers SYNOPSIS asz8[-pcpu][-llistfile][-ohexfile][-d][-ssymbolfile]input DESCRIPTION Theasz8commandassemblestheinputfileintoatextoutputfilerepresentingtheprogrammemoryforamicrocomputer. Options -pcpuOverridetheinstructionsetselection. ValidValuesforCpu z8600,z8,Z8,UPC,upc,Z8090 -llistfile Outputafileformatedwiththememoryaddress,data,andsourceinputlines. -ohexfile Outputthememorydataimageinaformacceptedbymostpromprogrammers. -hhexfile Sameas−o. −dSavetheintermediatedatafile(seeFILE)andaborttheexecutionattheendofprocessing. -ssymbolfile Printthesymboltablevaluesandnames,oneperlineinthespecifiedfile. FILES /usr/tmp/frtXXXXXX SEEALSO FrankensteinCrossAssemblersUsersManual(filebase.doc)Appendixforasz8Frankenstein Assembler(fileasz8.doc) NOTES Thereisonlyoneinputfile. Theprogramcanselectwhichsubsetofinstructionsisaccepted.Theprogramfirstusesthename withwhichitisinvoked.Thisisscannedforasubstringwhichindicateswhichsettouse.The-p optionsoverridesthisselctionbyperformingthesamesubstringsearch.Finallytheinputfilecan selectwhichsubsettousewiththeCPUstatement. Thereshouldonlybeoneexecutablefilewithallofitsnameslinkedtoit. 1